# VFMM3D: Releasing the Potential of Image by Vision Foundation Model for Monocular 3D Object Detection

### v1

@@ Abstract
Due to its cost-effectiveness and widespread availability, monocular 3D object detection, which relies solely on a single camera during inference, holds significant importance across various applications, including autonomous driving and robotics. Nevertheless, directly predicting the coordinates of objects in 3D space from monocular images poses challenges. Therefore, an effective solution involves transforming monocular images into LiDAR-like representations and employing a LiDAR-based 3D object detector to predict the 3D coordinates of objects. The key step in this method is accurately converting the monocular image into a reliable point cloud form. In this paper, we present VFMM3D, an innovative approach that leverages the capabilities of Vision Foundation Models (VFMs) to accurately transform single-view images into LiDAR point cloud representations. VFMM3D utilizes the Segment Anything Model (SAM) and Depth Anything Model (DAM) to generate high-quality pseudo-LiDAR data enriched with rich foreground information. Specifically, the Depth Anything Model (DAM) is employed to generate dense depth maps. Subsequently, the Segment Anything Model (SAM) is utilized to differentiate foreground and background regions by predicting instance masks. These predicted instance masks and depth maps are then combined and projected into 3D space to generate pseudo-LiDAR points. Finally, any object detectors based on point clouds can be utilized to predict the 3D coordinates of objects. Comprehensive experiments are conducted on the challenging 3D object detection dataset KITTI. Our VFMM3D establishes a new state-of-the-art performance, surpassing the best-reported method NeurOCS-MLC by 1.49% and 2.43% average mAP on KITTI val set in terms of 3D and BEV, respectively. Additionally, experimental results demonstrate the generality of VFMM3D, showcasing its seamless integration into various LiDAR-based 3D object detectors.
@@ I Introduction
The rise of autonomous systems has brought about an era where perceiving and understanding the environment in three dimensions is not just advantageous but imperative. Among the array of sensing methods, monocular vision emerges as a compelling choice due to its simplicity and cost-effectiveness. However, the challenge persists in extracting 3D information from a single 2D image, posing a significant obstacle. Monocular 3D object detection, tasked with estimating the 3D bounding boxes of objects from monocular images, stands as a pivotal element in numerous applications, spanning advanced driver-assistance systems (ADAS), robotics, and virtual reality, among others.
Traditional monocular 3D object detection approaches rely on geometry-based methods or depth estimation techniques, which often suffer from a lack of accuracy and robustness, particularly in complex and dynamic scenes. The recent surge in deep learning has introduced new possibilities, with convolutional neural networks (CNNs) and recent vision transformers (ViTs) [1], revolutionizing the field of computer vision. These models pre-trained on vast datasets have demonstrated an uncanny ability to generalize and perform well across a multitude of tasks.
Figure 1: Comparison with state-of-art methods on KITTI val set easy level 3D and BEV car class.
Vision Foundation Models (VFMs), such as the Segment Anything Model (SAM) [2] and the Depth Anything Model (DAM) [3], have shown particular promise in the domain of monocular 3D object detection. SAM with its ability to generate precise segmentation masks, and DAM with its capability to estimate precise scene depth from monocular images, provide a rich set of depth features that can significantly enhance the performance of monocular 3D detection models. The combination of these VFMs offers a unique opportunity to address the inherent challenges in monocular 3D detection, such as the ambiguity of depth and the need for robust segmentation.
In this paper, we present VFMM3D, a novel framework that synergistically integrates SAM and DAM to precisely perform monocular 3D object detection. Our approach leverages the strengths of both models to generate pseudo-LiDAR data enriched with semantic information and accurate depth. Based on the generated high-quality pseudo-LiDAR data, state-of-the-art LiDAR-based 3D object detectors could be employed to perform 3D object detection.
@@
Through extensive experiments on the KITTI dataset, we demonstrate that VFMM3D outperforms existing state-of-the-art monocular 3D object detection methods across different difficulty levels. The superior performance of VFMM3D is attributed to its ability to extract and utilize detailed 3D spatial information from monocular images effectively by VFMs. Our work not only pushes the frontier of monocular 3D object detection but also provides a robust and generalizable solution that can be readily applied to real-world applications.
@@
  • Extensive experiments on KITTI dataset show our method achieves state-of-art results among existing monocular 3d object detection methods as shown in Fig. 3.
@@ II Related Work
Monocular 3D Object Detection. In the field of monocular 3D object detection, a variety of approaches have been developed to infer 3D information from a single 2D image. Deep3DBox [6] introduces the MultiBin method for orientation estimation and employs geometric constraints of 2D bounding boxes to derive 3D bounding boxes. MonoPair [7] is a typical geometry-based method that uses spatial relationships between objects to estimate 3D properties. MonoFlex [8] proposes a flexible framework that separates truncated objects and amalgamates multiple depth estimation approaches, including direct regression and geometric solutions from keypoints. MonoDLE [9] and PGD [10] apply deep learning to directly predict 3D bounding boxes or depth on instances from images. MonoRUn [11] learns dense correspondences and geometry in a self-supervised manner, with simple 3D bounding box annotations. MonoDTR [12] and MonoDETR [13] adaptive feature aggregation via a depth-guided transformer for monocular 3D object detection. Pseudo-LiDAR-based methods [14, 15, 16, 17], transform the image into a pseudo-LiDAR point cloud and process it using voxel-based techniques. CaDDN [18] constructs bird’s-eye-view (BEV) representations by learning categorical depth distributions for each pixel. It then recovers bounding boxes from the BEV projection. MonoPSR [19] estimates the instance point cloud and enforces alignment between object appearance and the projected point cloud for proposal refinement. FQNet [20] organizes pseudo-LiDAR data into image representation and utilizes potent 2D CNNs to enhance detection performance. Addressing the problem with a 3D region proposal network, M3D-RPN [21] proposes depth-aware convolutional layers to enrich extracted features. RTM3D [22] predicts the projected vertices of 3D bounding boxes and resolves 3D properties through nonlinear least squares optimization. In this paper, we adopt a method based on pseudo-LiDAR, considering its superior generalizability, which enables seamless integration with various detectors to accommodate diverse scenarios.
@@ III-F LiDAR-Based Detection
Specifically, we demonstrate that VFMM3D works with three different LiDAR detectors: PointPillars [30], PV-RCNN [32], and Voxel-RCNN [31]. These are all widely-used LiDAR detectors with distinct network architectures: single-stage (PointPillars) versus two-stage (PV-RCNN and Voxel-RCNN).
Voxel-RCNN. The Voxel-RCNN method is designed to enhance the performance of voxel-based 3D object detection while maintaining computational efficiency. It introduces a two-stage framework that includes a 3D backbone network for feature extraction, a 2D bird-eye-view (BEV) Region Proposal Network (RPN), and a detection head. A key innovation is the voxel RoI pooling, which directly extracts RoI features from voxel features for further refinement. This approach allows for real-time frame processing while achieving comparable detection accuracy with state-of-the-art point-based models, significantly reducing computation costs.
TABLE I: Comparsion of our model with state-of-the-art models on the KITTI val set for the car class. ’Mod.’ indicates the moderate difficulty level. Red indicates the best results, while blue indicates the second-best results.
@@
NeurOCS-MLC [47] CVPR 2023 31.24 21.01 17.70 39.16 26.78 23.63 - - - - - -
VFMM3D(PointPillars) - 27.23 17.15 13.95 38.02 24.93 20.79 69.94 48.75 42.98 74.12 54.37 48.09
VFMM3D(Voxel-RCNN) - 34.60 21.58 18.23 44.18 28.66 24.02 72.85 51.83 45.53 76.69 55.43 48.95
VFMM3D(PV-RCNN) - 32.06 21.00 17.49 43.30 28.88 24.72 73.21 53.51 46.88 77.09 58.37 51.69
@@
PointPillars. PointPillars is a method designed for efficient object detection from point clouds, particularly for autonomous driving applications. It introduces a novel encoder that uses PointNets to learn a representation of point clouds organized in vertical columns (pillars). The encoded features can be used with any standard 2D convolutional detection architecture, and PointPillars proposes a lean downstream network for this purpose. By operating on pillars instead of voxels, PointPillars avoids the need for hand-tuning the binning of the vertical direction and leverages the efficiency of 2D convolutions on a GPU. The method significantly outperforms previous encoders in both speed and accuracy, running at 62 Hz and setting new standards for performance on KITTI benchmarks for both 3D and bird’s eye view detection.
@@ IV-A Settings
Dataset. We evaluate our method on the KITTI 3D object detection and bird’s eye view benchmark [48], which comprises 7,481 training samples and 7,518 testing samples, along with the corresponding LiDAR point clouds, stereo images, and full camera matrix. We divided the original training samples into a train spilt set with 3,712 samples and a validation split set with 3,769 samples following the previous methods [14, 16]. It is worth noting that during both training and testing phases, our approach exclusively utilizes single-view RGB image data and does not incorporate any LiDAR point cloud or stereo image data.
@@
Evaluation Metrics. We report the performance of our method with the standard evaluation metric $AP_{40}$ [49], in terms of average precision sampled at 40 recall positions in the precision-recall curve, under three different difficulties (easy, moderate, and hard). For broad validation, we selected two IoU thresholds for car class on 3D boxes and BEV boxes, namely 0.7 ($AP@0.7$) and 0.5 ($AP@0.5$).
@@
Figure 4: Qualitative results of VFMM3D on KITTI val set. We visualize our 3D bounding box estimates (in red) alongside ground truth annotations (in green) on front view images (1st and 3rd rows) and pseudo-LiDAR point clouds (2nd and 4th rows).
@@
Figure 5: Visualization of segmentation results (2nd row) from text-prompt SAM, depth maps (3rd row) from DAM, rwa Pseudo-LiDAR (4th row) from depth map and Painted Sparse Pseudo-LiDAR (5th row) results from each component of VFMM3D on KITTI.
@@
Training Details. The size of images is set to 1224$\times$370 pixels without any 2D data augmentation. In the LiDAR-based detection stage, we use 3D random horizontal flip augmentation and point-shuffle augmentation. Due to the inaccuracy of depth estimation, there is a significant deviation between the pseudo points inside the 3D annotated boxes and the real points. Therefore, in this method, we do not utilize the data augmentation method of GT-sampling [29]. For main experiments as in Tab. I and Tab. II, we use Voxel-RCNN [31] as the LiDAR-based 3D object detector. For Voxel-RCNN and PV-RCNN [32], we set the pseudo-LiDAR range as [0, 70.4], [-40.0, 40.0], [-3.0, 1.0] meters along the X, Y, and Z axes, respectively. And the voxel size is (0.05, 0.05, 0.1) meters. For PointPillars [30], we set the pseudo-LiDAR range as [0, 69.12], [-39.68, 39.68], [-3.0, 1.0] meters, the pillar size is set as (0.16, 0.16, 4) meters.
@@
Our method is primarily implemented using OpenPCDet [50]. We utilize pre-trained models, including Grounding DINO, SAM with vit-h, and DAM with vit-l, which are fine-tuned on KITTI and NYUv2 outdoor metric depth datasets. The LiDAR-based 3D object detectors are trained on the sparse painted pseudo-LiDAR using 4 RTX 3090 GPUs. We employ the AdamW [51] optimizer with a learning rate of 1e-4 and weight decay of 0.01. The batch size is set to 32 for PointPillars, 32 for Voxel-RCNN, and 8 for PV-RCNN. Additionally, the number of training epochs is set to 80 for all detectors.
IV-B Comparison with SOTA methods

### v2

@@ Abstract
Due to its cost-effectiveness and widespread availability, monocular 3D object detection, which relies solely on a single camera during inference, holds significant importance across various applications, including autonomous driving and robotics. Nevertheless, directly predicting the coordinates of objects in 3D space from monocular images poses challenges. Therefore, an effective solution involves transforming monocular images into LiDAR-like representations and employing a LiDAR-based 3D object detector to predict the 3D coordinates of objects. The key step in this method is accurately converting the monocular image into a reliable point cloud form. In this paper, we present VFMM3D, an innovative framework that leverages the capabilities of Vision Foundation Models (VFMs) to accurately transform single-view images into LiDAR point cloud representations. VFMM3D utilizes the Segment Anything Model (SAM) and Depth Anything Model (DAM) to generate high-quality pseudo-LiDAR data enriched with rich foreground information. Specifically, the Depth Anything Model (DAM) is employed to generate dense depth maps. Subsequently, the Segment Anything Model (SAM) is utilized to differentiate foreground and background regions by predicting instance masks. These predicted instance masks and depth maps are then combined and projected into 3D space to generate pseudo-LiDAR points. Finally, any object detectors based on point clouds can be utilized to predict the 3D coordinates of objects. Comprehensive experiments are conducted on two challenging 3D object detection datasets, KITTI and Waymo. Our VFMM3D establishes a new state-of-the-art performance on both datasets. Additionally, experimental results demonstrate the generality of VFMM3D, showcasing its seamless integration into various LiDAR-based 3D object detectors.
@@ I Introduction
The rise of autonomous systems has brought about an era where perceiving and understanding the environment in three dimensions is not just advantageous but imperative. Among the array of sensing methods, monocular vision emerges as a compelling choice due to its simplicity and cost-effectiveness. However, the challenge persists in extracting 3D information from a single 2D image, posing a significant obstacle. Monocular 3D object detection [1, 2, 3, 4, 5], tasked with estimating the 3D bounding boxes of objects from monocular images, stands as a pivotal element in numerous applications, spanning advanced driver-assistance systems (ADAS), robotics, and virtual reality, among others.
Traditional monocular 3D object detection approaches rely on geometry-based methods [6, 7, 8, 9, 10, 11] or depth estimation techniques [12, 13, 14, 15, 16, 17, 18], which often suffer from a lack of accuracy and robustness, particularly in complex and dynamic scenes. The recent surge in deep learning has introduced new possibilities, with convolutional neural networks (CNNs) and vision transformers (ViTs) [19], revolutionizing the computer vision field. These models, pre-trained on extensive datasets, have demonstrated an uncanny ability to generalize and perform well across a multitude of tasks.
Vision Foundation Models (VFMs), such as the Depth Anything Model (DAM) [20] and the Segment Anything Model (SAM) [21], have shown particular promise in the domain of monocular 3D object detection. SAM with its ability to generate precise segmentation masks, and DAM with its capability to estimate precise scene depth from monocular images, provide a rich set of depth features that can significantly enhance the performance of monocular 3D detection models. The combination of these VFMs offers a unique opportunity to address the inherent challenges in monocular 3D detection, such as the ambiguity of depth and the need for robust segmentation.
In this paper, we present VFMM3D, a novel framework that synergistically integrates SAM and DAM to precisely perform monocular 3D object detection. Our approach leverages the strengths of both models to generate pseudo-LiDAR data enriched with semantic information and accurate depth. The ability of our framework may be improved further by more powerful vision foundation models and LiDAR-based 3D object detectors in the future. Based on the generated high-quality pseudo-LiDAR data, state-of-the-art LiDAR-based 3D object detectors could be employed to perform 3D object detection.
@@
Through extensive experiments on the KITTI dataset, we demonstrate that VFMM3D surpasses existing state-of-the-art monocular 3D object detection methods across different difficulty levels. The superior performance of VFMM3D is attributed to its ability to effectively extract and utilize detailed 3D spatial information from monocular images by VFMs. Our work not only pushes the frontier of monocular 3D object detection but also provides a robust and generalizable solution that can be readily applied to real-world applications.
@@
  • Extensive experiments on KITTI and Waymo datasets show our method achieves state-of-art results among existing monocular 3d object detection methods as shown in Fig. 2.
@@ II Related Work
Monocular 3D Object Detection. In the field of monocular 3D object detection, a variety of methods have been developed to extract 3D information from a single 2D image. Deep3DBox [8] introduces the MultiBin approach for yaw estimation and utilizes geometric constraints of 2D bounding boxes to generate 3D bounding boxes. MonoPair [6] is a typical geometry-based method that uses spatial relationships between objects to estimate 3D properties. MonoFlex [24] proposes a flexible framework that separates truncated objects and amalgamates multiple depth estimation approaches, including direct regression and geometric solutions from keypoints. MonoDLE [16] and PGD [25] apply deep learning to directly predict 3D bounding boxes or depth on instances from images. MonoRUn [26] leverages self-supervised learning to establish dense correspondences and geometric relationships by using 3D bounding box annotations directly. MonoDTR [27] and MonoDETR [28] adaptive feature aggregation via a depth-guided transformer for monocular 3D object detection. Pseudo-LiDAR-based methods [29, 30, 31, 32] generate a pseudo-LiDAR point cloud from the images and then process it using voxel-based techniques. CaDDN [17] constructs bird’s-eye-view (BEV) representations by learning categorical depth distributions for each pixel. It then recovers bounding boxes from the BEV projection. MonoPSR [7] estimates instance point clouds and refines proposals by enforcing alignment between object appearance and the projected point cloud. QNet [33] transforms pseudo-LiDAR data into an image representation and employs powerful 2D CNNs to improve detection performance. To tackle this issue, M3D-RPN [12] introduces depth-aware convolutional layers within a 3D region proposal network to enhance the extracted features. RTM3D [34] estimates the projected vertices of 3D bounding boxes and resolves 3D properties through nonlinear least squares optimization. In this paper, we adopt a pseudo-LiDAR-based framework, considering its superior generalizability, which enables seamless integration with various detectors to accommodate diverse scenarios.
@@ III-F LiDAR-Based Detection
Specifically, we show that VFMM3D is compatible with three different LiDAR-based detectors: PV-RCNN [44], Voxel-RCNN [43] and PointPillars [42]. These include widely-used LiDAR detectors, each featuring a distinct network architecture: single-stage (PointPillars) versus two-stage (PV-RCNN and Voxel-RCNN).
Voxel-RCNN. The Voxel-RCNN method is designed to enhance the performance of voxel-based 3D object detection while maintaining computational efficiency. It introduces a two-stage framework that includes a 3D backbone network for feature extraction, a 2D bird-eye-view (BEV) Region Proposal Network (RPN), and a detection head. The key innovation is voxel RoI pooling, which straightforwardly extracts RoI features from voxel features for subsequent refinement. This approach allows for real-time frame processing while achieving comparable detection accuracy to that of state-of-the-art point-based approaches, significantly reducing computation costs.
TABLE I: Comparsion Result on Waymo val. set for the vehicles class. Red indicates the best results, while blue indicates the second-best results.
Diff. Extra Method Reference 3D mAP / mAPH(IoU=0.7) 3D mAP / mAPH(IoU=0.5)
Overall 0 - 30m 30 - 50m 50m - $\infty$ Overall 0 - 30m 30 - 50m 50m - $\infty$
L1 LiDAR CaDDN [17] CVPR 21 5.03 / 4.99 14.54 / 14.43 1.47 / 1.45 0.10 / 0.10 17.54 / 17.31 45.00 / 44.46 9.24 / 9.11 0.64 / 0.62
LiDAR MonoNeRD [53] ICCV 23 10.66 / 10.56 27.84 / 27.57 5.40 / 5.36 0.72 / 0.71 31.18 / 30.70 61.11 / 60.28 26.08 / 25.71 6.60 / 6.47
LiDAR DID-M3D [54] ECCV 22 - / - - / - - / - - / - 20.66 / 20.47 40.92 / 40.60 15.63 / 15.48 5.35 / 5.24
Depth PatchNet [55] ECCV 20 0.39 / 0.37 1.67 / 1.63 0.13 / 0.12 0.03 / 0.03 2.92 / 2.74 10.03 / 9.75 1.09 / 0.96 0.23 / 0.18
Depth PCT [56] NeurIPS 21 0.89 / 0.88 3.18 / 3.15 0.27 / 0.27 0.07 / 0.07 4.20 / 4.15 14.70 / 14.54 1.78 / 1.75 0.39 / 0.39
- M3D-RPN [12] ICCV 19 0.35 / 0.34 1.12 / 1.10 0.18 / 0.18 0.02 / 0.02 3.79 / 3.63 11.14 / 10.70 2.16 / 2.09 0.26 / 0.21
- GUPNet [9] ICCV 21 2.28 / 2.27 6.15 / 6.11 0.81 / 0.80 0.03 / 0.03 10.02 / 9.94 24.78 / 24.59 4.84 / 4.78 0.22 / 0.22
- DEVIANT [57] ECCV 22 2.69 / 2.67 6.95 / 6.90 0.99 / 0.98 0.02 / 0.02 10.98 / 10.89 26.85 / 26.64 5.13 / 5.08 0.18 / 0.18
- MonoJSG [58] CVPR 22 0.97 / 0.95 4.65 / 4.59 0.55 / 0.53 0.10 / 0.09 5.65 / 5.47 20.86 / 20.26 3.91 / 3.79 0.97 / 0.92
- SSD-MonoDETR [59] TIV 2023 4.54 / - 9.93 / - 1.18 / - 0.15 / - 11.83 / - 27.69 / - 5.33 / - 0.85 / -
- VFMM3D(Voxel-RCNN) - 7.48 / 7.24 17.41 / 16.91 2.08 / 1.99 0.28 / 0.24 18.26 / 17.46 38.26 / 36.75 8.35 / 7.95 1.60 / 1.36
- VFMM3D(PV-RCNN) - 8.06 / 7.79 19.38 / 18.81 2.22 / 2.05 0.15 / 0.12 18.39 / 17.48 39.26 / 37.66 8.81 / 8.06 1.10 / 0.87
L2 LiDAR CaDDN CVPR 21 4.49 / 4.45 14.50 / 14.38 1.42 / 1.41 0.09 / 0.09 16.51 / 16.28 44.87 / 44.33 8.99 / 8.86 0.58 / 0.55
LiDAR MonoNeRD [53] ICCV 23 10.03 / 9.93 27.75 / 27.48 5.25 / 5.21 0.60 / 0.59 29.29 / 28.84 60.91 / 60.08 25.36 / 25.00 5.77 / 5.66
LiDAR DID-M3D [54] ECCV 22 - / - - / - - / - - / - 19.37 / 19.19 40.77 / 40.46 15.18 / 15.04 4.69 / 4.59
Depth PatchNet [55] ECCV 20 0.38 / 0.36 1.67 / 1.63 0.13 / 0.11 0.03 / 0.03 2.42 / 2.28 10.01 / 9.73 1.07 / 0.94 0.22 / 0.16
Depth PCT [56] NeurIPS 21 0.66 / 0.66 3.18 / 3.15 0.27 / 0.26 0.07 / 0.07 4.03 / 3.99 14.67 / 14.51 1.74 / 1.71 0.36 / 0.35
- M3D-RPN [12] ICCV 19 0.33 / 0.33 1.12 / 1.10 0.18 / 0.17 0.02 / 0.02 3.61 / 3.46 11.12 / 10.67 2.12 / 2.04 0.24 / 0.20
- GUPNet [9] ICCV 21 2.14 / 2.12 6.13 / 6.08 0.78 / 0.77 0.02 / 0.02 9.39 / 9.31 24.69 / 24.50 4.67 / 4.62 0.19 / 0.19
- DEVIANT [57] ECCV 22 2.52 / 2.50 6.93 / 6.87 0.95 / 0.94 0.02 / 0.02 10.29 / 10.20 26.75 / 26.54 4.95 / 4.90 0.16 / 0.16
- MonoJSG [58] CVPR 22 0.91 / 0.89 4.64 / 4.65 0.55 / 0.53 0.09 / 0.09 5.34 / 5.17 20.79 / 20.19 3.79 / 3.67 0.85 / 0.82
- SSD-MonoDETR [59] TIV 23 4.12 / - 8.87 / - 1.02 / - 0.13 / - 11.34 / - 27.62 / - 5.21 / - 0.76 / -
- VFMM3D(Voxel-RCNN) - 6.62 / 6.41 17.17 / 16.67 1.89 / 1.81 0.22 / 0.19 16.19 / 15.48 37.77 / 36.29 7.58 / 7.21 1.25 / 1.06
- VFMM3D(PV-RCNN) - 7.14 / 6.89 19.11 / 18.55 2.01 / 1.86 0.12 / 0.09 16.31 / 15.50 38.78 / 37.19 8.01 / 7.33 0.86 / 0.68
TABLE II: Comparsion of our model with state-of-the-art models on the KITTI val. set for the car class. ’Mod.’ indicates the moderate difficulty level. * means the training set of model’s depth estimator may overlap the validation set. Red indicates the best results, while blue indicates the second-best results.
@@
PDR [4] TCSVT 2023 27.65 19.44 16.24 35.59 25.72 21.35 - - - - - -
VFMM3D(PointPillars) - 22.43 15.45 13.92 35.85 24.88 22.75 63.52 45.60 41.79 68.26 48.62 44.75
VFMM3D(Voxel-RCNN) - 29.09 19.41 17.09 39.43 26.56 23.69 68.72 50.53 45.02 73.17 52.90 47.25
VFMM3D(PV-RCNN) - 29.05 19.10 16.86 41.78 28.53 25.61 69.95 51.78 47.01 74.24 54.98 49.69
VFMM3D(PointPillars)* - 27.23 17.15 13.95 38.02 24.93 20.79 69.94 48.75 42.98 74.12 54.37 48.09
VFMM3D(Voxel-RCNN)* - 34.60 21.58 18.23 44.18 28.66 24.02 72.85 51.83 45.53 76.69 55.43 48.95
VFMM3D(PV-RCNN)* - 32.06 21.00 17.49 43.30 28.88 24.72 73.21 53.51 46.88 77.09 58.37 51.69
@@
PointPillars. PointPillars is a method designed for efficient object detection from point clouds, particularly for autonomous driving applications. It introduces a novel encoder that employs PointNets to learn representations of point clouds organized into vertical columns or pillars. The encoded features can be integrated with any standard 2D convolutional detection architecture, and PointPillars introduces a streamlined downstream network for this integration. By using pillars rather than voxels, PointPillars eliminates the need for hand-tuning the binning of the vertical direction and leverages the efficiency of 2D convolutions on a GPU. The method significantly outperforms previous encoders in both speed and accuracy, running at 62 Hz and setting new standards for performance on KITTI benchmarks for both 3D and bird’s eye view detection.
@@ IV-A Settings
Datasets. We evaluate our method on two datasets, Waymo and KITTI. The Waymo dataset [64] contains 1,150 video sequences collected from diverse driving environments. We use the official splitting protocol to split the dataset into a train set with 798 sequences, 158081 samples and a validation set with 202 sequences, 39,848 samples. We follow DEVIANT [57], PCT [56] and other methods [53, 58], using the front view camera for monocular 3D object detection. The KITTI dataset contains two benchmarks, 3D object detection and bird’s eye viewk [65], which comprises 7,481 training samples and 7,518 testing samples, along with the corresponding LiDAR point clouds, stereo images, and full camera matrix. We divided the original training samples into a train spilt set with 3,712 samples and a validation split set with 3,769 samples following the previous methods [29, 31]. It is worth noting that during both training and testing phases, our approach exclusively utilizes single-view RGB image data and does not incorporate any LiDAR point cloud or stereo image data.
@@
Evaluation Metrics. On Waymo, we follow the official evaluation metrics, evaluate on two object levels: Level 1 and Level 2 with $mAP$ and $mAPH$. It assigns each object to a level according to the number of LiDAR points contained within its 3D bounding box. Besides, we also provide the performance on three distance ranges: [0m, 30m), [30m, 50m), [50m, $\infty$]. For KITTI, we report the performance of our method with the standard evaluation metric $AP_{40}$ [66], in terms of average precision sampled at 40 recall positions in the precision-recall curve, under three different difficulties (easy, moderate, and hard). For broad validation, we selected two IoU thresholds for car class on 3D boxes and BEV boxes, namely 0.7 ($AP@0.7$) and 0.5 ($AP@0.5$).
@@
Figure 3: Qualitative results of VFMM3D on KITTI val. set. We visualize our 3D bounding box estimates (in red) alongside ground truth annotations (in green) on front view images (1st and 3rd rows) and pseudo-LiDAR point clouds (2nd and 4th rows).
@@
Figure 4: Visualization of segmentation results (2nd row) from text-prompt SAM, depth maps (3rd row) from DAM, raw Pseudo-LiDAR (4th row) from depth map and Painted Sparse Pseudo-LiDAR (5th row) results from each component of VFMM3D on KITTI.
@@
Training Details. The size of images is set to 1224$\times$370 pixels on the KITTI dataset and 1920$\times$ 1280 pixels on the Waymo dataset without any 2D data augmentation. In the LiDAR-based detection stage, we use 3D random horizontal flip augmentation and point-shuffle augmentation. Due to the inaccuracy of depth estimation, there is a significant deviation between the pseudo points inside the 3D annotated boxes and the real points. Therefore, in this method, we do not utilize the data augmentation method of GT-sampling [41]. We use Voxel-RCNN [43] as the LiDAR-based 3D object detector for experiments on the KITTI dataset. For Voxel-RCNN and PV-RCNN [44], we set the pseudo-LiDAR range as [0, 70.4], [-40.0, 40.0], [-3.0, 1.0] meters along the X, Y, and Z axes, respectively. And the voxel size is (0.05, 0.05, 0.1) meters. For PointPillars [42], we set the pseudo-LiDAR range as [0, 69.12], [-39.68, 39.68], [-3.0, 1.0] meters, the pillar size is set as (0.16, 0.16, 4) meters. On the Waymo dataset, we use Voxel-RCNN and PV-RCNN and set the pseudo-LiDAR range as [0, 75.2], [-75.2, 75.2], [-2.0, 4.0] meters with voxel size (0.1, 0.1, 0.15) meters and [0, 59.6], [-25.6, 25.6], [-2.0, 4.0] meters with voxel size (0.05, 0.05, 0.15) meters, respectively.
@@
Our method is primarily implemented using OpenPCDet [67]. We utilize pre-trained models, including Grounding DINO, SAM with vit-h, and DAM with vit-l, which is fine-tuned on KITTI outdoor metric depth datasets for containing metric depth estimation ability. However, we note the officially provided DAM model is finetuned on the whole KITTI-depth dataset, as mentioned in DD3D [30] and Pesuo-Lidar++ [50], this training set overlap with the KITTI-3D validation data for detection. Thus, we use the KITTI-3D clean training split, which removes training images that are geographically close to any of the KITTI-3D images provided by DD3D to finetune the DAM to avoid any bias. We also demonstrate the results on KITTI with DAM finetuned with the whole KITTI-depth dataset as some methods [60] also report this kind of results, as shown in Tbale II with the indicator *. The LiDAR-based 3D object detectors are trained on the sparse painted pseudo-LiDAR using 8 RTX A6000 GPUs. We employ the AdamW [68] optimizer with a learning rate of 4e-4 and weight decay of 0.01. The batch size is set to 32 for PointPillars, 32 for Voxel-RCNN, and 8 for PV-RCNN. Additionally, the number of training epochs is set to 80 for all detectors.
IV-B Waymo Results
Tab. I shows the performance comparison results on Waymo validation spilt. It demonstrates our method VFMM3D outperforms other monocular methods without extra modality data during training [12, 9, 57, 58], which surpasses the latest monocular method DEVIANT [57] by 7.41% and 6.20% on level 1 and level 2 3D mAP(IoU=0.5), respectively. It is worth noting that our method even outperforms CaDNN [17], which uses LiDAR as extra training data, 0.72% on level 1 3D mAP(IoU=0.5).
IV-C KITTI Results